OpenStudy (anonymous):

the cost of fabric varies directly with the amount of fabric purchased. if 1 2/3 yards of fabric cost 10.00 how much would 4 1/2 yards cost?

OpenStudy (anonymous):

Varies directly means y = kx you solve for k and then use it in what you are trying to find.

OpenStudy (anonymous):

cf = cost of fabric fp = fabric purchased. cf = k times fp or y = kx

OpenStudy (anonymous):

10.00 = k(1 2/3) 1 2/3 is 5/3 10.00 = k(5/3) multiply both sides by 3/5 3/5(10.00) = k(5/3)(3/5) 6 = k

OpenStudy (anonymous):

Now you start over. y = kx k is 6 so y = 6x

OpenStudy (anonymous):

You are trying to find the cost which is y y = 6(4.5) 4 1/2 yards. y = $27.00

OpenStudy (anonymous):

oh my gosh thankkkkkkkkk you so much :) that helped a lot!